# Exploring Pessimism and Optimism Dynamics in Deep Reinforcement Learning

### v1

@@ Appendix B USAC under gaussianity
Proof of Proposition 2.
The proof follows by similar steps as in the proof of Proposition 1.
Similarly, we can construct a USAC algorithm under Gaussian uncertainty, analogous to our approach under Laplace uncertainty as discussed in Section 4.3. By using the utility function defined in 16, we can incorporate specific $\lambda_{\text{critic}}$ and $\lambda_{\text{actor}}$ parameters for critic and actor training, respectively. This allows us to adjust the levels of optimism and pessimism tailored to the Gaussian-distributed $\mathcal{Q}$ values, thereby facilitating effective learning and exploration in our framework.
Appendix C Experiments details
Table 3: Properties of MuJoCo environments
Environment Action dimension Observation dimension
Ant-v4 8 27
HalfCheetah-v4 6 17
Hopper-v4 3 11
Humanoid-v4 17 376
Walker2d-v4 6 17
Environments
Our experiment is implemented in PyTorch [21, Version $2.1.0$]. The experiments are conducted on 5 continuous control environments in the MuJoCo physics engine [30, 5]. More detailed information about these environments are provided in Table 3.
Table 4: Shared hyper-parameters
Hyper-parameter Value
@@ Appendix C Experiments details
Replay ratio 1
@@
Evaluation and seeds
As mentioned in Table 4, we train our algorithms for one million time steps. Each experiment is repeated across 5 different seeds. We evaluate at every $10,000$ time step for ten evaluation episodes. For training, we use seeds 1 to 5, and for evaluation, we use seeds 101 to 105, which correspond to the training repetition number plus 100.
Hyper-parameters
The hyper-parameters and network configurations used in our experiments can be found in Table 4. For SAC and USAC, we learn the entropy temperature $\alpha$ during as outlined in Haarnoja et al. [13]. For OAC [7] and TOP [20], we adopt the best parameters reported in their respective papers.
Architectures and baselines
The architectural specifications for all algorithms are presented in Table 5. Here, $d_{s}$ and $d_{a}$ represent the dimensions of the state and action spaces, respectively. The SquashedGaussian module features a Gaussian head, which uses the first $d_{a}$ inputs for mean calculation and the subsequent $d_{a}$ inputs for variance. This architecture is the standard one used in original implementations of SAC [12], TD3 [10], and OAC [7]. However, while TOP [20] states that it employs a network architecture with $2$ hidden layers, its original implementation code actually considers $3$ hidden layers. [^6] We use the implementation provided by the authors for TOP, https://github.com/tedmoskovitz/TOP, where the "number of random actions" at the start of training updated to $10000$. This discrepancy prompted us to re-run TOP with a synchronized network such that it aligns with the other algorithms. In addition, the critic network in TOP follows the original implementation with a return shape that equals the number of quantiles (N-Quantiles), which is $50$.
Table 5: Architecture details
Actor network Critic network
USAC
SAC
OAC
TD3
TOP
USAC
SAC
OAC
TD3
TOP
Linear($d_{s}$, 256)
ReLU()
Linear($d_{s}$ + $d_{a}$, 256)
ReLU()
Linear(256, 256)
ReLU()
Linear(256, 256)
ReLU()
Linear(256, 2×$\times d_{a}$, $\theta$)
Linear(256, $d_{a}$) Linear(256, 1) Linear(256, N-Quantiles)
SquashedGaussian(2×$\times d_{a}$, $d_{a}$) Tanh()
Grid search of USAC
In order to explore the dynamics of pessimism and optimism in USAC with Laplace distribution (Section 4.3), we conduct a grid search across five continuous control environments in MuJoCo (see e.g., environments specifications above). These grid search results are averaged over 3 seeds (seed 1 to 3), each with 3 evaluation episodes. This grid search investigated various values of $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$, using SAC as the baseline for comparison.
We examined critics ranging from pessimistic to semi-optimistic and actors from pessimistic to optimistic. Specifically, we used $\kappa_{\text{critic}}\in\{-0.831559,-0.5,-0.33\}$ [^7] A $\kappa=-0.831559$ corresponds to $g(\kappa)=1$, which means that the utility function in 11 simplifies into the minimum of the two critics. Hence, by setting $\kappa_{\text{critic}}\approx-0.83$, we align with the pessimistic choices of SAC [12], TD3 [10], and OAC [7]., and $\kappa_{\text{actor}}\in\{-0.99,-0.5,0,0.5,0.99\}$. The rationale for choosing more pessimistic (to semi-optimistic) $\kappa_{\text{critic}}$ values is to counteract the overestimation bias. Given that USAC stems from the framework of SAC, SAC was included in our figures for comparison.
In Figures 2, 3, and 4, we present the results of our grid search, showcasing the final returns, estimation error, and area under the learning curves for various parameter combinations. The estimation error calculated by subtracting the estimated $Q$-value from the true (discounted) reward. Thus, positive values indicate underestimating, while negative value indicate overestimation. These findings provide insights into the dynamics of pessimism and optimism in our USAC algorithm. As mentioned in Section 5, they reveal that: (i) overestimation (negative estimation error) should always be avoided as it reduces performance, and (ii) underestimation (positive estimation error) does not necessarily harm performance. This suggests that a pessimistic approach should be adopted in critic training, while either a pessimistic or optimistic approach can be suitable for actor training, depending on the environment. The best combinations of $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ in terms of final returns are summarized in Table 2. These optimal pairs highlight the importance of tailoring the levels of pessimism and optimism to the specific requirements of each environment to enhance overall performance.
@@
Figure 4: Grid search results for the area under the learning curve of USAC algorithm with Laplace distribution, averaged over three seeds, each with three evaluation episodes. The grid search explore various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
Learning curve and estimation error curves
The learning curve and estimation error curves of our USAC algorithm and the baseline algorithms can be seen in Figures 5 and 6, which correspond to the experiments summarized in Table 1. In addition, the estimation error values for these experiments are provided inTable 6. For USAC, the legend corresponds to the default case where $\kappa_{\text{critic}}=\kappa_{\text{actor}}=-0.831559$. The USAC (best) legend uses the optimal $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ pairs (with respect to final return) as listed in Table 2.For the Humanoid-v4 environment, we found that the default parameter pair outperformed the best pair identified through grid search, e.g., see Figure 2. Therefore, we used the default pair for this environment in our results (Table 1). However, we conjecture that by expanding our (limited) grid search, a better pair may exist.
Table 6: Estimation error of USAC and baselines.
@@
Default $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ Best $(\kappa_{\text{critic}},\kappa_{\text{actor}})$
Ant-v4 $80.66$ $91.84$ $75.33$ $196.77$ $205.49$ $98.99$
HalfCheetah-v4 $221.21$ $146.07$ $222.93$ $216.54$ $190.28$ $81.97$
Hopper-v4 $12.46$ $18.51$ $12.73$ $24.82$ $25.05$ $11.73$
Humanoid-v4 $40.24$ $40.24^{\star}$ $42.78$ $31.89$ $52.66$ $24.02$
Walker2d-v4 $42.50$ $44.87$ $36.65$ $50.88$ $52.55$ $34.34$
@@
Figure 6: Estimation error curves of USAC and baselines. Solid lines indicate the estimated value function by the critic, while dashed lines show the averaged episodic discounted true return. The shaded areas around the solid and dashed lines represent the standard deviation from the average estimated value and the average true reward, respectively.
Computational cost
Our USAC algorithm uses twin critics and a single actor network as for SAC, TD3, OAC, and TOP. This design choice eliminates the need for any additional computational steps and is no reliance on repetitive sampling, pre-training, or iterative approximation procedures. Consequently, it can be conclude that its computational characteristics closely match those of other widely used actor-critic algorithms and ours comparing baselines reported. The code for our algorithm implementation is also available alongside the paper.

### v2

@@ 4.3 Laplace distributed critics
Proof of Proposition˜3.
The proof follows by similar steps as in the proof of Proposition˜2. ∎
Now suppose the uncertainty follows a Laplace distribution. Following the algorithmic steps outlined in Section˜4.2, we can estimate the utility of the target critic distribution $\bar{\mathcal{Q}}$ by two target critics $\{\bar{Q}_{k}\}_{k=1,2}$, which are delayed versions of the critics $\{Q_{k}\}_{k=1,2}$. The utility function is given by $U^{\bar{\mathcal{Q}}}_{\kappa_{\text{critic}}}=\mu_{\bar{\mathcal{Q}}}+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}$ with utility parameter $\kappa_{\text{critic}}\in(-1,1)$, where
$\mu_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a)),$ (14)
and
$\sigma_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}\lvert\bar{Q}_{1}(s,a)-\bar{Q}_{2}(s,a)\rvert$. (15)
Hence, the target value $y(s,a,r,s^{\prime})$ to train the critic becomes
$r+\gamma[\mu_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],$ (16)
with $a^{\prime}\sim\pi(\cdot|s^{\prime})$. Similarly, the actor learns its policy by
$\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[\mu_{\tilde{\mathcal{Q}}}(s,a)+g(\kappa_{\text{actor}})\sigma_{\tilde{\mathcal{Q}}}(s,a)-\alpha\log\pi(a|s)],$ (17)
for $\kappa_{\text{actor}}\in(-1,1)$, where $\mu_{\tilde{\mathcal{Q}}}$ and $\sigma_{\tilde{\mathcal{Q}}}$ are estimated in similar manner as ˜14 and 15 on $\tilde{\mathcal{Q}}$ using the critics $\{Q_{k}\}_{k=1,2}$. The pseudo-code of this algorithm can be found in Algorithm˜1.
We note that a smaller $\kappa$ value indicates more pessimism, while a larger $\kappa$ value indicates more optimism. By choosing $\kappa$ such that $g(\kappa)=-1$ (i.e., $\kappa\approx-0.83$), the utility in ˜11 simply reduces to the minimum of the two critics. Thus, choosing $\kappa_{\text{critic}}=\kappa_{\text{actor}}\approx-0.83$, we obtain the pessimism choices of SAC [14] and TD3 [12]. On the other hand, by choosing $\kappa_{\text{critic}}\approx-0.83$ and a positive $\kappa_{\text{actor}}$, we have a pessimism critic and an optimistic actor as explored in OAC [9]. Lastly, by letting a multi-armed bandit algorithm choose $\kappa_{\text{critic}}=\kappa_{\text{actor}}\in\{-0.916563,0\}$, we recover the pessimistic/optimistic design suggested by TOP [22]. To recover the pessimistic case of the TOP algorithm, we can choose $\kappa=-0.916563$, which corresponds to having the function $g$ in ˜12 equal to $\sqrt{2}$. We note that by choosing 0.916563-0.916563, TOP considers a more pessimistic critic target than SAC, TD3, and OAC.
5 Experiments
We conduct experiments to answer two questions: (Q1) Can our utility-theoretic framework explain the previously observed relationships between the action-value biases and exploration optimism by two simple scalars, one for actor and one for the critic? (Q2) Can these scalars be steered to improve performance via a coarse grid-search?
Throughout the experiments, we adopted the USAC version described in Algorithm˜1 that assumes Laplace-distributed critics as the representative implementation of our framework. We tested all models on the standard five continuous control benchmarks of the MuJoCo physics engine [33, 7] to facilitate the interpretation of our results and its comparability to previous studies. The implementation was done in PyTorch [23, Version $2.1.0$]. We ran each experiment for one million time steps and tested each policy every $10,000$ steps over ten evaluation episodes. We repeated each experiment across five different seeds hard-set to numbers 1 to 5 for training. We chose seed numbers for evaluation as the training repetition number plus 100. We consider SAC [14], TD3 [12], OAC [9], and TOP [22] as baselines, as they represent the most powerful algorithms that take finding an optimal balance between optimism and pessimism as the main goal.  [^1] For TOP, we use https://github.com/tedmoskovitz/TOP with the number of random actions set to 10,000.
We modeled all the actor and critic networks as fully connected layers of $256$ hidden units that follow ReLU() activations. USAC, SAC, and OAC actors output parameters for a SquashedGaussian module via a linear layer of size $2\times d_{a}$, where the first outputs represent the mean and the second represent the variance. TD3 and TOP use deterministic policies with a linear layer of size $d_{a}$ followed by Tanh(). Critic networks for USAC, SAC, OAC, and TD3 output scalars; TOP outputs $N=50$ quantiles. Here, $d_{s}$ and $d_{a}$ denote state and action dimensions. Note that although the paper of TOP describes a two-layer network, its code uses three layers; we modify it for consistency with other baselines. For SAC and USAC, we employ automatic entropy temperature tuning $\alpha$ [15]. For OAC and TOP, we adopt the best parameters reported in their respective papers. The hyperparameters and network configurations shared by all models in our consideration are summarized in Table˜1. The baseline models do not have any additional hyperparameters required to be tuned. We provide a public implementation at https://github.com/adinlab/UtilitySoftActorCritic.
Table 1: Shared hyperparameters
Hyperparameter Value
@@ 5 Experiments
Replay ratio/update-to-data 1
@@
Our framework can be implemented as a drop-in into any standard actor-critic pipeline as an alternative way to aggregate critic outputs into a single value. Hence, it does not cause a tangible computational overhead. Our representative USAC implementation adopts the twin-critic and single-actor architecture used in standard SAC, TD3, OAC, and TOP implementations. Its computation time is indistinguishable from a standard SAC implementation.
@@
Figure 2: The impact of the degree of actor optimism on the final return of USAC algorithm when Laplace distributed critics are assumed. The results are averaged over three seeds, each with three evaluation episodes. The experiments span various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
Figure 3: The impact of the degree of actor optimism on the estimation error (true-estimation) of USAC algorithm when Laplace distributed critics are assumed. The results are averaged over three seeds, each with three evaluation episodes. The experiments span various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
Figure 4: The impact of the degree of actor optimism on the learning speed of USAC algorithm when Laplace distributed critics are assumed, measured by the area under learning curve. The results are averaged over three seeds, each with three evaluation episodes. The experiments span various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
5.1 Question 1: Can USAC hyperparameters explain the optimism and pessimism dynamics?
We investigate the performance profile of USAC with Laplace distributed critics (Section˜4.3) across different degrees of optimism and pessimism on five MuJoCo environments. We expect to replicate the findings of prior art [12, 9, 22], thereby verifying our claim that the trade-off problem can be reduced to tuning an interpretable hyperparameter. Results are averaged over three different seeds, comparing against SAC as a baseline, as it is a special case of USAC with $\kappa_{\text{critic}}=\kappa_{\text{actor}}=-0.831559$. This allows us to systematically analyze how different configurations of utility parameters affect learning behavior and overall returns. We examined critics ranging from pessimistic to semi-optimistic and actors from pessimistic to optimistic. Specifically, we used $\kappa_{\text{critic}}\in\{-0.831559,-0.5,-0.33\}$ [^2] Recall that $\kappa=-0.831559$ corresponds to $g(\kappa)=1$, simplifying the utility function in 11 to the minimum of the two critics. Therefore, by setting $\kappa_{\text{critic}}\approx-0.83$, we align with the pessimistic choices of SAC [14], TD3 [12], and OAC [9]., and $\kappa_{\text{actor}}\in\{-0.99,-0.5,0,0.5,0.99\}$}. The rationale for choosing more pessimistic (to semi-optimistic) $\kappa_{\text{critic}}$ values is to counteract the overestimation bias commonly seen in off-policy reinforcement learning. We present results for three metrics that provide complementary information: (i) final return (Figure˜2) to assess the success of each configuration at solving the task at hand, (ii) estimation error (Figure˜3) to assess the explanatory value of our framework for further investigation, (iii) area under the learning curve (Figure˜4) to assess the learning speed of each configuration.
The estimation error is calculated by subtracting the estimated $Q$-value from the true (discounted) reward. Thus, positive values indicate underestimating, while negative values indicate overestimation, both of which can impact learning efficiency and stability. Our results verify our hypothesis that the following key findings of the prior art can indeed be explained by a single hyperparameter value:
  • Overestimation (i.e., negative estimation error) should always be avoided as it reduces performance, as visible from the strong correlation between high estimation error in Figure˜3 and low final return in Figure˜2.
  • Underestimation (i.e., positive estimation error) does not necessarily harm performance. This outcome reaffirms the central claims reported in the TD3 paper [12], which yielded the wide adoption of the min-clipping method.
We also conclude from the results that a pessimistic approach should be adopted in critic training, whereas in actor training, either a pessimistic or optimistic approach can be suitable depending on the environment and the exploration-exploitation balance required. This outcome reaffirms the prior findings [12, 9, 22], supporting our hypothesis about the possibility of expressing the optimism and pessimism dynamics on a simple and interpretable axis.
Table 2: The best pair of hyperparameters $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ for USAC with respect to final returns; see for example Figure˜2.
Environment $\kappa_{\text{critic}}$ $\kappa_{\text{actor}}$
Ant-v4 0.831559-0.831559 0.99-0.99
HalfCheetah-v4 0.33-0.33 0.50-0.50
Hopper-v4 0.831559-0.831559 $0.50$
Humanoid-v4 0.831559-0.831559 0.831559-0.831559
Walker2d-v4 0.831559-0.831559 $0.0$
5.2 Question 2: Can the hyperpararameters be effortlessly tuned to improve performance?
From Figures˜2, 3, and 4, we can extract the best combinations of $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ per environment with respect to final return and treat them as the outcome of a search on a coarse grid of hyperparameters. These best values are listed in Table˜2. The difference between the best values across the environments highlights the environment-specific nature of optimism and pessimism and also demonstrates the success of our framework at detecting these differences.
Table 3: Final return on MuJoCo environments trained with 1M time steps, averaged over 5 seeds. The best algorithms are highlighted in bold. $\pm$ corresponds to the standard deviation across repetitions. $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ for USAC are listed in Table˜2.
@@ 5.2 Question 2: Can the hyperpararameters be effortlessly tuned to improve performance?
Ant-v4 $\boldsymbol{5158\pm 424}$ $4756\pm 773$ $4091\pm 130$ $4177\pm 802$ $4334\pm 1276$
HalfCheetah-v4 $\boldsymbol{11736\pm 297}$ $10763\pm 891$ $10570\pm 801$ $8684\pm 1671$ $7311\pm 3074$
Hopper-v4 $\boldsymbol{3442\pm 126}$ $3185\pm 380$ $1986\pm 1114$ $293\pm 112 33$$369\pm 154$
Humanoid-v4 $\boldsymbol{5602\pm 210}$ $5503\pm 198$ $5149\pm 221$ $5390\pm 121$ $5332\pm 445$
Walker2d-v4 $\boldsymbol{4530\pm 708}$ $3757\pm 1069$ $4369\pm 594$ $3467\pm 1193$ $4317\pm 631$
@@
We give a performance comparison of the grid-optimized USAC against the baseline models in Table˜3. The superior performance of USAC supports our hypothesis that the hyperparameters introduced by our framework also have an instrumental value, as they can be used to improve actor-critic algorithms. It is also remarkable that USAC incurs a smaller standard deviation than the baseline models across repetitions in the far majority of the cases, indicating an improved model stability during learning. The learning curve of the grid-optimized USAC algorithm is given in Figure˜5 together with the baseline algorithm. The results suggest a consistent performance improvement that spans the whole learning regime.
In applications where online learning with minimum regret is more important than maximizing the final reward, such as interactive learning applications, a similar grid-search may also be performed with respect to the area under curve. While we omit a detailed reporting of such an analysis for brevity, one can see from Figure˜4 that USAC will be similarly competitive in that scenario.
6 Discussion and limitations
We introduced a novel algorithmic framework to trade pessimistic policy evaluation and pessimistic/optimistic policy search in off-policy deep actor-critic pipelines. The key idea is to balance optimism and pessimism through incorporating a utility function that captures the uncertainty in the critic due to limited access to the environment into a single and interpretable hyperparameter. This USAC framework is flexible and can be adapted to different distributions, with our experiments primarily focusing on the Laplace distribution.
Our USAC framework allows independent control over the optimism and pessimism levels for the actor and critic. This decoupling enables a stable critic, necessary for accurate value function estimation, and an explorative actor, crucial for effective policy learning. Furthermore, USAC allows tuning the degree of optimism/pessimism via a single hyperparameter that interpretably encodes full pessimism as -1 and full optimism as +1 and all intermediate possibilities as values in between. Our experiments revealed that being pessimistic in critic training while being either optimistic or pessimistic in actor training can be beneficial. However, it is also essential to consider the specific environment. We also found that a search on a coarse grid of our new hyperparameters brings consistent performance gains across environments with diverse characteristics.
The primary limitation of USAC is that it does not prescribe any insights for on-the-fly tuning of its hyperparameters. An exciting direction for future research is to develop adaptive schemes where these parameters adjust dynamically throughout the learning process to better respond to the changing dynamics of the environment. There are several potential approaches to integrate automatic tuning. For example, (i) adapting the automatic tuning scheme used for $\alpha$ in SAC [14, 15], (ii) employing bandit-based selection methods [22], or (iii) exploring continuous learning approaches using gradient descent [6].
Another limitation is our evaluation scope. We focus on the five most used MuJoCo environments [33], where the studied phenomenon and the impact of our solution are clearly visible, ensuring comparability with the key baselines we consider. Broader validation across additional domains would further demonstrate generality, which we plan to explore in future work.
Our findings could also be improved by incorporating higher moments of critic distributions (mean, variance, skewness, kurtosis) to better capture uncertainty in value estimates. Lastly, extending from critic pairs to ensembles may offer faster training at the cost of increased computational power.
Acknowledgements